ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 147, 2005

Alon, G. Cohomology of local systems coming from p-adic hyperplane arrangements

189

Božikov, Z. and Z. Janko On a question of N. Blackburn about finite 2-groups

329

Chorny, B. Localization with respect to a class of maps I — Equivariant localization of diagrams of spaces

93

Chorny, B. Localization with respect to a class of maps II — Equivariant cellularization and its application

141

Einsiedler, M. and T. Ward Isomorphism rigidity in entropy rank two

269

Green, B. and I. Z. Ruzsa Sum-free sets in abelian groups  157

Greschonig, G. Recurrence in unipotent groups and ergodic nonabelian group extensions

245

Janko, Z. See Z. Božnikov

Krueger, J. Destroying stationary sets  285

Lancien, G. and B. Randrianantoanina On the extension of Hölder maps with values in spaces of continuous functions

75

Randrianantoanina, B. See G. Lancien

Roblin, T. Un théorème de Fatou pour les densités conformes avec applications aux revêtements galoisiens en courbure négative

333

Ruzsa, I. Z. See B. Green

Schramm, O. Compositions of random transpositions  221

Sela, Z. Diophantine geometry over groups III: Rigid and solid solutions

1

Sodin, M. and B. Tsirelson Random complex zeroes, III. Decay of the hole probability

371

Tsirelson, B. See M. Sodin

Velickovic, B. CCC forcing and splitting reals  209

Ward, T. See M. Einsiedler

Yang, L. Z. The growth of linear differential equations and their applications 359